# Advances in 4D Representation: Geometry, Motion, and Interaction

### v1

@@ 5.3 Interaction Dataset
Human-Object Interaction. HOI datasets capture the coupled dynamics between humans and manipulated objects during interaction sequences. GRAB [TGBT20] provides grasping sequences with rigid objects, while BEHAVE [BXP22] extends to deformable object interactions using RGB-D sensors. Recent datasets like HUMOTO [LHB25] and ParaHome [KKNJ25] offer larger-scale captures with natural home environment interactions. These datasets enable learning of contact-rich interactions and object affordances, though capture complexity and annotation requirements limit dataset scale and diversity.
@@ 5.4 Benchmarks and Evaluation Metrics
Specialized Benchmarks. Recent benchmarks target specific 4D challenges: WideRange4D [YZT25] focuses on wide-range spatial movements of objects in synthetic scenes, Inter3D [CHY25] benchmarks on human-interactable 3D object reconstruction, SEED4D [KGB25] provides hours of annotated driving videos with syncrhonised ego-exo captures for autonomous driving benchmark; InterAct [XLZ25] provides  22 hours of HOI data for 6 HOI task benchmarks; 4D-Bench provides a benchmark on MLLM understanding on 4D by assessing the captioning abilities. Despite progress, unified benchmarks for core generation tasks (text-to-4D, image-to-4D, dynamic NeRF reconstruction) remain absent. Current practice relies on small-scale datasets like Consistent4D [JZG24] and D-NeRF [PCPMMN21] without standardized protocols.
@@
Quantitative Metrics. Evaluation metrics span multiple objectives. Appearance fidelity uses PSNR and SSIM for pixel-level correspondence, complemented by perceptually-aligned LPIPS [ZIE18] that better correlates with human quality judgments. Temporal consistency employs Fréchet Video Distance (FVD) commonly applied to 4D video renders [JZG24, ZJZ24], STREAM [KKY24] for disentangled evaluation of temporal coherence, visual fidelity, and diversity without length constraints, and VBench [HHY24] providing comprehensive assessment across 16 dimensions including subject consistency and motion smoothness. Semantic alignment measures input-output consistency through CLIP-score [RKH21] for text-to-image tasks and R-Precision for text-to-motion alignment via retrieval accuracy. Geometric integrity employs Chamfer Distance for bidirectional point cloud similarity, Earth Mover’s Distance (EMD) [RTG98] for optimal transport cost between point distributions, 3D IoU for volumetric overlap, Hausdorff Distance for maximum deviation, and physics-based constraints for interaction validity [LHB25]. Human evaluation complements automated metrics, assessing faithfulness, aesthetic quality, and physical plausibility through structured user studies [NCOZMA25, ZYW24].
@@
Current Limitations. Contemporary 4D evaluation faces fundamental challenges. Geometric functionality assessment remains inadequate, with metrics emphasizing shape alignment while overlooking functional correctness, structural balance, and design complexity. Long-duration motion evaluation lacks appropriate protocols, as temporal metrics are constrained to short sequences (<32 frames) without assessment capabilities for extended temporal consistency or motion degradation. Unified benchmark scarcity limits systematic progress assessment—the field lacks comprehensive benchmarks encompassing diverse object categories, motion types, and conditioning modalities comparable to ImageNet [DDS09]. Perceptual alignment gaps between quantitative scores and human perception necessitate heavy reliance on manual evaluation, motivating efforts toward automated perception-aligned metrics [WZM24]. Addressing these requires functionally-aware metrics, extended temporal assessment, comprehensive benchmarks, and improved perceptual alignment.
@@ 6 Training Strategy
The training strategies employed in 4D tasks can be broadly categorized into per-scene optimization, feed-forward models, and hybrid optimization, each reflecting a trade-off between generalization, fidelity, and efficiency (Figure˜6). Building upon the geometric representations (Section˜2) and motion modeling approaches (Section˜3) discussed above, we provide a comprehensive taxonomy of selected papers organized by these three dimensions in  Table˜1.
@@ 6.1.1 Data-driven Priors
Foundation model priors provide implicit knowledge from large-scale datasets as substitutes for scarce 4D supervision. Multi-modal diffusion priors address complementary quality aspects: image diffusion [RBL22] for appearance detail, 3D-aware image diffusion [SCZ23, LGL23] for spatial consistency, video diffusion [BDK23, SPH22, GYR23] for temporal dynamics, and multi-view diffusion [SWY24, LLZ23] for $360^{\circ}$ consistency. 4D-fy [BSR24] alternates supervision across all three types, achieving 67% user preference over MAV3D [SSP23]. Recent work utilizes MLLMs/VLMs for 4D generation [ZL25] and LLM reasoning for object composition [XLB24].
Auxiliary priors such as depth offers geometric cue, optical flow offers motion cues and semantic cues provide enforce region- or part-level consistency across time,thus stabilizing optimization, improving spatial-temporal coherence, and enhancing both geometric fidelity and perceptual realism in 4D tasks. Depth supervision improves joint learning of radiance and motion fields [CFF22, WKS22], particularly in outdoor scenes [WLL23, TLH24, YCW23]. Monocular depth estimation [RLH20, YKH24] provides geometric cues despite scale ambiguity [WYG24, LLW24a]. Optical flow models [XZC23] supply dense inter-frame correspondences [YVN22, LCM22], while visual foundation models like DINO [ODM24] deliver semantic features consistent across views and time [YJB23, WLJ23]. SA4D [JWF24] lifts SAM[KMR23]-style 2D masks into 4D Gaussian representations by learning a temporal identity feature field. In dynamic scenes, object-level semantic segmentation provides valuable silhouettes or foreground masks that help localize dynamic objects and decompose scenes into static backgrounds and dynamic foregrounds [YLL23, WLZ24, LGXC24]. For object articulation [GXL25], instead of using foundational model priors, it uses physics based optimization [IMH05] to learn the motion parameters.
@@ 6.1.2 Design Choices
The process of converting diverse 2D or textual inputs into 3D (and subsequently 4D) structure depends heavily on how the input modality is encoded and lifted. Here we discuss some of the design choices that has influenced the 4D literature in the realm of per-scene optimization.
3D lifting models differ in how they convert 2D/textual inputs into 3D structure. Multi-view generation approaches like Zero-1-to-3 [LWVH23] and MVDream [SWY24] for objects, as well as ViewCrafter [YXY24] and SEVA [ZGV25] for scenes, synthesize novel views and reconstruct structure through NeRF or Gaussian Splatting. This has recently been improved by multi-view video generators [VYB24] for better geometry quality. Direct feedforward methods [XLX25, TCC24, JN23] infer 3D structure from single images via learned mappings. Multi-view synthesis ensures stronger geometric consistency; feedforward prediction offers faster, generalizable inference at reduced fidelity cost.
@@
Deformation fields model motion patterns using MLPs [BKY24, HSY24, LKLR24], spatial-temporal planes [DMY23b, LDZ24, WYF24], polynomial functions [LCLX24], Fourier series [KVN25], or combinations [LDZY24b]. These methods learn motion patterns from independent time inputs in a vanilla way, neglecting internal cross-time relationships. Recent work [JKZ24] adaptively learns temporal relationships using transformer attention. Most methods optimize deformation from multi-view videos frame-by-frame [YXV25, ZJZ24], while articulation-based methods [LJL25, GXL25] estimate motion using heuristics or input geometry.
@@ 6.1.3 Optimization Objectives
where $\mathcal{T}$ is the number of video frame, $f(.)$ is the rendering function and $o_{\textrm{Ref}}$ is the reference camera. For object centric scenes [JZG24], a separate foreground masking loss [WYG24, ZJZ24] is also estimated These reconstruction losses complement or replace SDS when ground truth video supervision is available. These losses mostly enhance the per-frame appearance, reduces floating artifacts.
@@
Open challenges. Optimization-based 4D methods achieve high visual fidelity but remain impractical for large-scale use. The per-scene optimization is computationally expensive, unstable, and slow to converge, limiting scalability and motion complexity in real-world applications. Here we discuss some of the pitfalls.
Compute cost These methods are extremely time-consuming, with [ZJZ24] and [JZG24] requiring around 2 hours per scene and [SSP23] or [BSR24] taking several to $10+$ hours. This makes them 10100010-1000x slower than feedforward models, preventing real-time or interactive use. They also demand high-end GPUs like V100/A100 and large memory budgets (atleast 24 GB GPU RAM), posing significant hardware barriers.
Scalability Each scene must be optimized independently, eliminating the possibility of batch or parallel generation. Such slow, scene-specific pipelines are impractical for large-scale or commercial deployment, where generating hundreds of 4D assets for gaming, film, or AR/VR would be prohibitively costly. Moreover, learned representations do not generalize across scenes.
Limited motion complexity Current pipelines perform well for simple, object-scale motions but fail to capture large, scene-level dynamics, non-rigid deformations, or topological changes. Realistic physics is rarely enforced, and long-range motion across extended sequences remains unsolved. Modeling the appearance or disappearance of objects over time remains particularly challenging for continuous representations.
@@ 6.2 End-to-End Training
Feedforward models learn global mappings from 2D observations to 4D representations using large pre-trained networks. Once trained, these models infer novel scenes in single forward passes, enabling real-time generation and cross-dataset scalability. They offer significant speed advantages (e.g., L4GM [RXM24] reconstructs in 0.3s, 4DGT [XLD25] at 25 ms/frame), learn strong spatiotemporal priors handling ambiguity and sparse views with in-domain generalization, and provide deterministic inference with stable quality and fixed runtime regardless of scene complexity.
@@ 6.2.1 Data-driven Priors
Motion priors leverage video diffusion transformers for joint motion-appearance learning. 4DNex [CLZ25] integrates RGB and 3D sequences using Wan2.1 [WWA25] DiT for 6D video modeling. SV4D [YXV25] builds on Stable Video Diffusion for multi-view sequence synthesis, with v2.0 adding 3D-aware attention. DynamiCrafter-based methods [JZL25, YCPH25] extend image-to-video diffusion to output explicit geometry such as point maps, disparity, and Gaussian splats. SP4D [ZYD25] enhances SV4D with dual RGB/Part branches and BiDiFuse exchange for kinematic consistency, while DeepVerse [CZH25a] uses flow matching to generalize diffusion prior for long-horizon coherence. Hybrid pipelines like CAT4D [WGP25] combine CAT3D [GHH24] image-to-3D diffusion with Lumiere [BTCT24] video diffusion, and DreamArt [LLT25] fine-tunes SVD [BDK23] for articulated motion. Overall, diffusion-based models replace costly optimization with learned motion and view priors, achieving scalable and photorealistic 4D generation.
Geometry priors anchor dynamic learning in pre-trained reconstruction models. L4GM [RXM24], AR4D [ZHY25], and CAT4D [WGP25] extend text-to-3D pipelines using ImageDream [WS23], MVDream [SWY24], and CAT3D [GHH24]. Large 3D backbones provide spatial priors: L4GM builds on LGM [TCC24], 4D-LRM [MCY25] extends LRM [HZG23], Splat4D [YCPH25] fine-tunes LGM encoders, and Forge4D [HHC25] adopts VGGT. MonST3R [ZHH24] and Geo4D [JZL25] employ DUSt3R/CroCo encoders, while DepthAnythingV2 [YKH24], StableNormal [YQG24], LiDAR, and DeepLabv3+ [CZP18] refine geometric cues [XLD25, LXZ24]. 2D foundation models [KMR23, ODM24] enhance segmentation consistency [LLT25, ZYD25]. Human-object interaction methods [WHZ25b, LSCK23, LSN24] derive priors from motion capture datasets [LHB25]. Together, these priors transfer pre-trained 3D spatial understanding to dynamic 4D tasks, achieving high-quality results with minimal 4D supervision.
@@ 6.2.2 Design Choices
The fundamental design of a 4D model is defined by how it represents 3D space and how it captures the evolution of that space over time.
3D lifting model Most recent works adopt 3D Gaussian Splatting (3DGS) [KKLD23] for its balance of realism, efficiency, and differentiability. Two main extensions adapt it to 4D:
@@
Sequential or Deformable 3DGS: Other approaches construct per-frame 3DGS scenes linked by interpolation [RXM24, LRM24] or continuous deformation fields, as in CAT4D [WGP25] and AR4D [ZHY25], which evolve a canonical scene over time.
Beyond GS, mesh-based models (e.g., AnimateAnyMesh [WYWB25], DreamArt [LLT25]) handle articulated structures explicitly, while diffusion-based pipelines like SV4D [YXV25] rely on implicit NeRF-style fields [PSB21]. Overall, current trends favor explicit, geometry-aware, and temporally coherent 3D lifting frameworks that unify appearance and motion in a feedforward manner.
Motion models have evolved beyond simple interpolation. Explicit motion prediction methods [LLP25, HHC25] use transformer heads to predict dense 3D motion fields or per-pixel displacements. Temporal transformers [RXM24, YXV25] incorporate attention for cross-frame dependencies, enabling implicit learning of complex interactions—a popular choice in human-object interaction [LSN24, XSW23], recently replaced by diffusion methods [RTSB25, SZL24]. Autoregressive approaches [ZHY25, CZH25a] generate scenes sequentially for long-term coherence, while MonST3R [ZHH24] combines feedforward pairwise modeling with global temporal optimization. Together, these paradigms transition from per-frame fitting to learned feedforward or autoregressive abstractions, improving tem- poral continuity and scalability.

### v2

@@ 5.3 Interaction Dataset
Human-Object Interaction. HOI datasets capture the coupled dynamics between humans and manipulated objects during interaction sequences, spanning multiple levels of granularity. At the finest level, hand-object interaction datasets including DexYCB [CYX21], ARCTIC [FTT23], OAKINK2 [ZYZ24], and TACO [LYS24] capture single-hand or bi-manual grasping and manipulation with rigid objects, while Obman [HVT19] provides synthetic data pairing MANO [RTB17] hand models with ShapeNet [CFG15] objects. Whole-body interaction datasets extend this scope, with CIRCLE [ALV23] capturing reaching motions in virtual home environments, COUCH [ZBS22] and CHAIRS [JLC23] focusing on sitting-based furniture interactions, and GRAB [TGBT20], BEHAVE [BXP22], and OMOMO [LWL23] providing diverse activities such as grasping, carrying, and moving household objects. At the broadest level, Human-Scene Interaction (HSI) datasets such as ParaHome [KKNJ25], SAMP [HCV21], TRUMANS [JZL24], and HUMANISE [WCL22] incorporate environmental navigation by capturing humans moving through spaces while interacting with multiple scene objects. These datasets enable learning of contact-rich interactions and object affordances, though capture complexity and annotation requirements limit dataset scale and diversity. Given the breadth of this field, the datasets presented here are not exhaustive; we refer readers to [ZMR23] for a comprehensive discussion of HOI datasets.
@@ 5.4 Benchmarks and Evaluation Metrics
Specialized Benchmarks. Recent benchmarks target specific 4D challenges: WideRange4D [YZT25] focuses on wide-range spatial object movements in synthetic scenes, Inter3D [CHY25] benchmarks human-interactable 3D object reconstruction, SEED4D [KGB25] provides annotated driving videos with synchronized ego-exo captures for autonomous driving tasks; InterAct [XLZ25] provides 22 hours of HOI data for 6 HOI task benchmarks; 4D-Bench [ZLZ25] benchmarks MLLM understanding on 4D by assessing captioning abilities. Despite progress, unified benchmarks for core generation tasks (text-to-4D, image-to-4D, dynamic NeRF reconstruction) remain absent. Current practice relies on small-scale datasets like Consistent4D [JZG24] and D-NeRF [PCPMMN21] without standardized protocols.
@@
Quantitative Metrics. Evaluation metrics span multiple objectives. Appearance fidelity uses PSNR and SSIM for pixel-level correspondence, complemented by perceptually-aligned LPIPS [ZIE18] that better correlates with human judgments. Temporal consistency employs Fréchet Video Distance (FVD) on 4D video renders [JZG24, ZJZ24], STREAM [KKY24] for disentangled evaluation of temporal coherence, visual fidelity, and diversity without length constraints, and VBench [HHY24] providing comprehensive assessment across 16 dimensions, including subject consistency and motion smoothness. Semantic alignment measures input-output consistency through CLIP-score [RKH21] for text-to-image tasks and R-Precision for text-to-motion alignment via retrieval accuracy. Geometric integrity employs Chamfer Distance for bidirectional point cloud similarity, Earth Mover’s Distance (EMD) [RTG98] for optimal transport cost between point distributions, 3D IoU for volumetric overlap, Hausdorff Distance for maximum deviation, and physics-based constraints for interaction validity [LHB25]. Human evaluation complements automated metrics, assessing faithfulness, aesthetic quality, and physical plausibility through structured user studies [NCOZMA25, ZYW24].
@@
Current Limitations. Contemporary 4D evaluation faces fundamental challenges. Geometric functionality assessment remains inadequate, with metrics emphasizing shape alignment while overlooking functional correctness, structural balance, and design complexity. Long-duration motion evaluation lacks appropriate protocols, as temporal metrics assess only short sequences (<32 frames) without extended temporal consistency evaluation. Unified benchmark scarcity limits systematic progress assessment—the field lacks comprehensive benchmarks encompassing diverse object categories, motion types, and conditioning modalities comparable to ImageNet [DDS09] on 2D tasks. Perceptual alignment gaps between quantitative scores and human perception necessitate heavy reliance on manual evaluation, motivating efforts toward automated perception-aligned metrics [WZM24]. Addressing these requires functionally-aware metrics, extended temporal assessment, comprehensive benchmarks, and improved perceptual alignment.
@@ 6 Training Strategy
The training strategies employed in 4D tasks reflect fundamental trade-offs between generalization, fidelity, and efficiency. For single-entity geometry and motion modeling, these strategies fall into three categories: per-scene optimization (Section˜6.1), feed-forward models (Section˜6.2), and hybrid approaches that combine both (Section˜6.3). Training interaction models, however, introduces distinct challenges, namely coordinating multi-entity dynamics under physical constraints, that warrant separate treatment (Section˜6.4). Building upon the geometric representations (Section˜2) and motion modeling approaches (Section˜3) discussed above, we provide a comprehensive taxonomy of selected papers organized by these dimensions in Table˜1.
@@ 6.1.1 Data-driven Priors
Foundation model priors provide implicit knowledge from large-scale datasets as substitutes for scarce 4D supervision. Multi-modal diffusion priors address complementary quality aspects: image diffusion [RBL22] for appearance detail, 3D-aware image diffusion [SCZ23, LGL23] for spatial consistency, video diffusion [BDK23, SPH23, GYR23] for temporal dynamics, and multi-view diffusion [SWY24, LLZ23] for $360^{\circ}$ consistency. 4D-fy [BSR24] alternates supervision across all three types, achieving 67% user preference over MAV3D [SSP23]. Recent work utilizes MLLMs/VLMs for 4D generation [ZL25] and LLM reasoning for object composition [XLB24].
Auxiliary priors such as depth offer geometric cues, optical flow offers motion cues, and semantic priors enforce region- or part-level consistency across time, thus stabilizing optimization, improving spatial-temporal coherence, and enhancing both geometric fidelity and perceptual realism in 4D tasks. Depth supervision improves joint learning of radiance and motion fields [CFF22, WKS22], particularly in outdoor scenes [WLL23, TLH24, YCW23]. Monocular depth estimation [RLH20, YKH24] provides geometric cues despite scale ambiguity [WYG24, LLW25]. Optical flow models [XZC23] supply dense inter-frame correspondences [YVN22, LCM22], while visual foundation models like DINO [ODM24] deliver semantic features consistent across views and time [YJB23, WLJ23]. SA4D [JWF24] lifts SAM [KMR23]-style 2D masks into 4D Gaussian representations by learning a temporal identity feature field. In dynamic scenes, object-level semantic segmentation provides valuable silhouettes or foreground masks that help localize dynamic objects and decompose scenes into static backgrounds and dynamic foregrounds [YLL23, LGXC24] . For object articulation [GXL25], instead of using foundational model priors, it uses physics-based optimization [IMH05] to learn motion parameters.
@@ 6.1.2 Design Choices
Converting 2D or textual inputs into 3D (and subsequently 4D) structure depends heavily on how the input modality is encoded and lifted. Here we discuss some of the design choices that have influenced 4D literature in the realm of per-scene optimization.
3D lifting models differ in how they convert 2D/textual inputs into 3D structure. Multi-view generation approaches like Zero-1-to-3 [LWVH23] and MVDream [SWY24] for objects, as well as ViewCrafter [YXY24] and SEVA [ZGV25] for scenes, synthesize novel views and reconstruct structure through NeRF or Gaussian Splatting. This has recently been improved by multi-view video generators [VYB24] for better geometry quality. Direct feed-forward methods [XLX25, TCC24, JN23] infer 3D structure from single images via learned mappings. Multi-view synthesis ensures stronger geometric consistency; feed-forward prediction offers faster, generalizable inference at reduced fidelity cost.
@@
Deformation fields model motion patterns using MLPs [BKY24, HSY24, LKLR24], spatial-temporal planes [DMY24, LDZ24, WYF24], polynomial functions [LCLX24], Fourier series [KVN24], or combinations [LDZY24]. These methods directly learn motion patterns from independent time inputs, neglecting cross-time relationships. Recent work [JKZ24] adaptively learns temporal relationships via transformer attention. Most methods optimize deformation frame-by-frame from multi-view videos  [YXV25, ZJZ24], while articulation-based methods [LJL25, GXL25] estimate motion from heuristics or input geometry.
@@ 6.1.3 Optimization Objectives
where $\mathcal{T}$ is the frame count, $f(.)$ is the rendering function and $o_{\textrm{Ref}}$ is the reference camera. For object centric scenes [JZG24], a separate foreground mask loss [WYG24, ZJZ24] is also estimated. These reconstruction losses complement or replace SDS when ground truth video supervision is available. These losses primarily enhance per-frame appearance and reduce floating artifacts.
@@
6.1.4 Open Challenges
Optimization-based 4D methods achieve high visual fidelity, but remain impractical for large-scale use due to computational expense, instability, and slow convergence, limiting scalability and motion complexity in real-world applications. Key limitations include:
Compute cost. These methods are extremely time-consuming, with [ZJZ24] and [JZG24] requiring around 2 hours per scene and [SSP23] or [BSR24] taking several to $10+$ hours. This makes them 10100010-1000x slower than feed-forward models, preventing real-time or interactive use. They also demand high-end GPUs like V100/A100 and large memory budgets (at least 24 GB GPU RAM), posing significant hardware barriers.
Scalability. Each scene must be optimized independently, eliminating the possibility of batch or parallel generation. Such slow, scene-specific pipelines are impractical for large-scale or commercial deployment, where generating hundreds of 4D assets for gaming, film, or AR/VR would be prohibitively costly. Moreover, learned representations do not generalize across scenes.
Limited motion complexity. Current pipelines perform well for simple, object-scale motions but fail to capture large, scene-level dynamics, non-rigid deformations, or topological changes. Realistic physics is rarely enforced, and long-range motion across extended sequences remains unsolved. Modeling the appearance or disappearance of objects over time remains particularly challenging for continuous representations.
@@ 6.2 End-to-End Training
Feed-forward models learn global mappings from 2D observations to 4D representations using large pre-trained networks. Once trained, these models infer novel scenes in single forward passes, enabling real-time generation (e.g., L4GM [RXM24] at 0.3s, 4DGT [XLD25] at 25 ms/frame) and cross-dataset scalability with learned spatiotemporal priors that handle ambiguity and sparse views, and provide deterministic inference with stable quality and fixed runtime regardless of scene complexity.
@@ 6.2.1 Data-driven Priors
Motion priors. Some methods leverage video diffusion models for joint motion-appearance learning. 4DNeX [CLZ25] integrates RGB and 3D sequences using DiT architectures. Geo4D [JZL25] extends image-to-video diffusion to output explicit geometry with point maps. Other works, such as AnimateAnyMesh [WYWB25] and Motion 3-to-4 [CCZ26] directly learn motion prior from large animated 3D assets.
Geometry priors. Large 3D backbones are widely adopted to anchor geometry priors: L4GM builds on LGM [TCC24], 4D-LRM [MCY25] extends LRM [HZG23], and Forge4D [HHC25] adopts VGGT. MonST3R [ZHH24] and Geo4D [JZL25] employ DUSt3R encoders. Human-object interaction methods [WHZ25, LSCK23, LSN24] anchor on predefined template geometries [PCG19, RTB17]. Together, these priors transfer pre-trained 3D spatial understanding to dynamic 4D tasks, achieving high-quality results with minimal 4D supervision.
@@ 6.2.2 Design Choices
The fundamental design of a 4D model is defined by how it represents 3D space and captures the evolution of that space over time.
3D lifting models. Most recent works adopt 3D Gaussian Splatting (3DGS) [KKLD23] for its balance of realism, efficiency, and differentiability. Two main extensions adapt it to 4D:
@@
Sequential or Deformable Fields: Other approaches construct sequential 3DGS scenes linked by interpolation [RXM24], continuous flow [LXZ24] or conditioned on timestamps [LRM24]. Beyond GS,  [WYWB25, CCZ26] handle mesh animations through vertex displacement maps. Overall, current trends favor explicit, geometry-aware, and temporally coherent 3D lifting frameworks that unify appearance and motion in a feed-forward manner.
Motion models have evolved beyond simple interpolation. Explicit motion prediction methods [LLP25, HHC25] use transformer heads to predict dense 3D motion fields or per-pixel displacements. Temporal transformers [RXM24] incorporate attention for cross-frame dependencies, enabling implicit learning of complex interactions—a popular choice in human-object interaction [LSN24, XSW23], recently replaced by diffusion methods [RTSB25, SZL24]. Together, these paradigms transition from per-frame fitting, in-between bridging to learned feed-forward representations, improving temporal continuity and scalability.